# The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders in the Randall-Sundrum model

### v5

@@ IV Conclusion
In this paper, by using Feynman diagram techniques we have evaluated the influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The results show that with fixed collision energies, the total cross-sections for hadronic productions in final state depend strongly on the parameters of the unparticle physics and muon beam polarizes. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficients as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$. For more details we have evaluated the statistical significance for the considered process. The numerical values indicate that the effect is greatly enhanced in the future experiments with moderately high energy colliders.

### v6

@@ IV Conclusion
In this paper, by using Feynman diagram techniques we have evaluated the influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The results show that with fixed collision energies, the total cross-sections for hadronic productions in final state depend strongly on the parameters of the unparticle physics, muon beam polarizes and the anomalous couplings. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficients as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$.
     For more details we have evaluated the statistical significance for the considered process. The numerical values indicate that with the contribution of the new physics from the RS model including unparticle and anomalous couplings at the experimental bound of the W branching ratio of $\mathcal{O}(10^{-6})$, the effect is much enhanced and can be measured at the multi-TeV muon collider. The best sensitivities on the anomalous coupling $WW\gamma$ can reach at ($10^{-1}$ - $10^{-2}$) which are much larger than that on the anomalous coupling $WWZ$ under the same conditions.